# Multi-user QKD using quotient graph states derived from continuous-variable dual-rail cluster states

### v1

@@ 7 Results
8 Conclusion
@@
Looking ahead, the insights gained from this work highlight several promising directions for future research. Exploring alternative multipartite states with configurations that allow users to access multiple modes could further enhance key rates and robustness in quantum communication protocols. Extending the protocol to networks with more than three users and examining the performance benefits in larger, more complex quantum networks is a natural progression. Developing advanced methods for channel parameter estimation that leverage access to multiple modes per user could further mitigate finite-size effects and improve the security and efficiency of the protocols. Pursuing experimental validation of the proposed protocol, including the generation of quotient graph states and testing under realistic conditions with impure squeezed states, would bridge the gap between theory and practice. Additionally, extending the security analysis to consider more sophisticated eavesdropping strategies would verify the protocol’s robustness against a wider range of attacks.
@@ Acknowledgements
I would like to acknowledge Vladyslav Usenko, Ivan Derkach, Tobias Gehring, Christoph Pacher, Florian Kanitschar, and Ulrik Lund Andersen for their valuable discussions. A. O. acknowledge the project 8C22002 (CVStar), which have received funding from the European Union’s Horizon 2020 research and innovation framework programme under grant agreement No. 731473 and 101017733, and the project No. 21-44815L of the Czech Science Foundation. A. O. acknowledges the project IGA-PrF-2024-008 of Palacky University Olomouc.
@@ Appendix A Dual rail cluster state to six mode graph state
In Sec.5, we introduced a method to generate a pure six-mode graph state from a dual-rail cluster state. Here, we provide a justification for this method. Since we are working with Gaussian states, it suffices to show that, after appropriately grouping the nodes as suggested by the coloring scheme in Fig.2(a), the covariance matrix of the resulting state corresponds to that of a pure six-mode graph state.
The dual-rail cluster state, denoted here by $\rho_{\text{dual}}$, consists of two parallel chains (rails) of quantum states: Rail $A$ and Rail $B$. Each node (quantum state) in these rails is uniquely identified by its rail ($A$ or $B$) and its position $i$ along that rail as illustrated bellow.
@@
where $\operatorname{Tr}^{(R)}_{i}$ denotes the partial trace over all modes except the one at position $i$ in rail $R$. We then form statistical mixtures by averaging the constituent states within each group. For each group $j\in\{1,2,3\}$ in rail $R$, the mixed state $\boldsymbol{\rho}^{(R)}_{j}$ is defined as:
$\boldsymbol{\rho}^{(R)}_{j}=\frac{1}{n}\sum_{m=1}^{n}\rho^{(R)}_{3m-(3-j)},$ (19)
@@
By forming these statistical mixtures, we create six mixed states in total (three groups in each of the two rails). The quadrature operators associated with these mixed states are denoted $\boldsymbol{x}^{(R)}_{j}$ and $\boldsymbol{p}^{(R)}_{j}$, where $j\in\{1,2,3\}$ and $R\in\{A,B\}$.
@@
$\left\langle\left(\boldsymbol{X}^{(R)}_{j}\right)^{2}\right\rangle=\frac{1}{n}\sum_{m=1}^{n}\left\langle\left(X^{(R)}_{3m-(3-j)}\right)^{2}\right\rangle,$ (20)
$\frac{1}{2}\left\langle[\boldsymbol{X}^{(R)}_{p},\boldsymbol{X}^{(S)}_{q}]_{+}\right\rangle=\frac{1}{2n}\sum_{m=1}^{n}\left\langle[X^{(R)}_{3m-(3-p)},X^{(S)}_{3m-(3-q)}]_{+}\right\rangle,$ (21)
@@
$\frac{1}{2}\left\langle[\boldsymbol{X}^{(k)}_{1},\boldsymbol{X}^{(l)}_{3}]_{+}\right\rangle=\frac{1}{2n}\sum_{m=2}^{n}\left\langle[X^{(k)}_{3m-3},X^{(l)}_{3m-2}]_{+}\right\rangle$ (22)
@@
$\mathbf{X}=\left(\boldsymbol{x}^{(A)}_{1},\boldsymbol{p}^{(A)}_{1},\boldsymbol{x}^{(B)}_{1},\boldsymbol{p}^{(B)}_{1},\boldsymbol{x}^{(A)}_{2},\boldsymbol{p}^{(A)}_{2},\boldsymbol{x}^{(B)}_{2},\boldsymbol{p}^{(B)}_{2},\boldsymbol{x}^{(A)}_{3},\boldsymbol{p}^{(A)}_{3},\boldsymbol{x}^{(B)}_{3},\boldsymbol{p}^{(B)}_{3}\right)^{T}.$ (23)
With this ordering, the covariance matrix $\boldsymbol{\sigma}$ of the six-mode graph state is given by:
$\boldsymbol{\sigma}=\begin{pmatrix}\boldsymbol{V}&\boldsymbol{0}&\boldsymbol{C}&-\boldsymbol{C}&\frac{n-1}{n}\boldsymbol{C}&\frac{n-1}{n}\boldsymbol{C}\\ \boldsymbol{0}&\boldsymbol{V}&\boldsymbol{C}&-\boldsymbol{C}&-\frac{n-1}{n}\boldsymbol{C}&-\frac{n-1}{n}\boldsymbol{C}\\ \boldsymbol{C}&\boldsymbol{C}&\boldsymbol{V}&\boldsymbol{0}&\boldsymbol{C}&-\boldsymbol{C}\\ -\boldsymbol{C}&-\boldsymbol{C}&\boldsymbol{0}&\boldsymbol{V}&\boldsymbol{C}&-\boldsymbol{C}\\ \frac{n-1}{n}\boldsymbol{C}&-\frac{n-1}{n}\boldsymbol{C}&\boldsymbol{C}&\boldsymbol{C}&\boldsymbol{V}&\boldsymbol{0}\\ \frac{n-1}{n}\boldsymbol{C}&-\frac{n-1}{n}\boldsymbol{C}&-\boldsymbol{C}&-\boldsymbol{C}&\boldsymbol{0}&\boldsymbol{V}\end{pmatrix}.$ (24)
Here, $\boldsymbol{V}$ and $\boldsymbol{C}$ are $2\times 2$ matrices representing the variances and covariances of the quadrature operators, respectively. These can be expressed as a function related to the squeezed quadrature variance $v$ and excess AS noise of the squeezed state employed in generating the dual rail cluster state Eq.17. Specifically, $\boldsymbol{V}$ is the covariance matrix of a single mixed state, and $\boldsymbol{C}$ captures the correlations between different mixed states.
The factor $\frac{n-1}{n}$ within the covariance matrix are due to the fact that the count of vertices in a group that includes the extreme vertices cannot match the number of edges linking them. this factor makes the multi partite Gaussian state impure. One can purify this impurity, introduced due to finite states, by introducing additional modes. Alternatively, one can consider pure six mode covariance matrix by removing the factor $\frac{n-1}{n}$. Introducing additional modes to purify the covariance matrix Eq.24 would give us a tighter bound than considering the covariance matrix Eq.15 . that is, considering the pure six mode covariance matrix to estimate Holevo bound is more pessimistic than considering a purified state corresponding to covariance matrix Eq.24, to estimate the Holevo bound. For sufficiently large $n$, the covariance matrix $\boldsymbol{\sigma}$ can be approximated to Eq.15 a pure six mode graph state and hence the improvement achieved through purification is insignificant.
@@ Appendix B Congruence Transformation for Covariance Matrices
Given an original covariance matrix $\boldsymbol{\Sigma}$ of dimensions $12n\times 12n$, of the modes of the dual-rail cluster state under consideration for grouping, we can derive the covariance matrix $\boldsymbol{\sigma}$ of dimensions $12\times 12$ corresponding to the six-mode graph state. This is achieved by applying a congruence transformation using a matrix $\mathbf{A}$ of size $12\times 12n$, such that:
$\boldsymbol{\sigma}=\frac{1}{n}\mathbf{A}\boldsymbol{\Sigma}\mathbf{A}^{\intercal}.$ (25)
@@ B.1 Congruence transformation for adjacency matrices
A complex-weighted adjacency matrix is used in continuous-variable quantum computing to represent Gaussian pure states, extending the real-weighted graphs of ideal cluster states to more realistic, physical states [31]. This formalism simplifies the analysis of entanglement structures, Gaussian operations, and deviations from ideal behaviors, which are essential for applications in CV quantum computing. Thus, we define a similar transformation to the one used in the previous section, enabling us to derive the adjacency matrix for a six-mode graph state $\boldsymbol{Z}_{six-mode}$ from the adjacency matrix of a dual-rail cluster state $\boldsymbol{Z}_{dual}$.
$\boldsymbol{Z}_{six-mode}=\frac{1}{n}\boldsymbol{B}\boldsymbol{Z}_{dual}\boldsymbol{B^{T}}$ (28)
@@
(b) Complex weighted graph of qoutent graph derived from dual rail cluster state.
@@
Appendix C Parameter estimation
@@
Here, $x^{(i)}$ denotes a squeezed quadrature with variance $v$, while $p^{(i)}$ denotes the corresponding anti-squeezed quadrature with variance $dv+\frac{1}{v}$.
@@
$\mathrm{Var}[x_{d}\tilde{x}_{u}]=\frac{t}{16}\left(5\left(dv+\frac{1}{v}\right)^{2}+5v^{2}+6\left(dv+\frac{1}{v}\right)v\right)+(1-t)\left(\frac{1+v(dv+v)}{2v}\right).$ (46)
@@
$x_{1}=\frac{\sqrt{2}\,p^{(1)}+x^{(2)}+\sqrt{3}\,x^{(3)}}{\sqrt{6}},\quad x_{2}=\frac{\sqrt{2}\,x^{(2)}-p^{(1)}}{\sqrt{3}},\quad x_{3}=\frac{-\sqrt{2}\,p^{(1)}-x^{(2)}+\sqrt{3}\,x^{(3)}}{\sqrt{6}}.$ (47)
@@
$\mathrm{Var}[x_{d}\tilde{x}_{u}]=\frac{t}{9}\left(2\left(dv+\frac{1}{v}\right)^{2}+5v^{2}+2\left(dv+\frac{1}{v}\right)v\right)+(1-t)\left(\frac{1+v(dv+2v)}{3v}\right).$ (48)
@@
$x_{d}=\frac{x^{(1)}+p^{(2)}}{\sqrt{2}},\quad x_{u}=\frac{1}{2}(\sqrt{2}\,p_{0}-x^{(1)}+p^{(2)}).$ (49)
@@
$\mathrm{Var}[x_{d}\tilde{x}_{u}]=\frac{t}{8}\left(dv+\frac{1}{v}\right)\left(dv+\frac{1}{v}+1\right)+v(1+v)+(1-t)\left(\frac{1+v(dv+v)}{2v}\right).$ (50)

### v2

@@ VII Results
Table 2: Recommended distribution strategies and corresponding information reconciliation methods for multi-user QKD
Strategy Operational Context Information Reconciliation Advantages
Distribution 1 Dealer is also a user Direct Reconciliation, Reverse Reconciliation Optimal for maximal distance; robust against trusted noise.
Distribution 2 Dealer is also a user Direct Reconciliation, Reverse Reconciliation Better performance in finite-size regime; enhanced channel parameter estimation via multiple correlations.
Distribution 3 The dealer distributes three modes. Measures and publicly discloses the remaining three. Entanglement-in-the-Middle Suitable for decentralized conference key agreement.
Distribution 4 Fully distributed; no dealer involvement after state distribution. Entanglement-in-the-Middle Enables fully peer-to-peer operation.
Distribution 5 Dealer sends two uncorrelated modes to remote users Direct Reconciliation (CKA), Reverse Reconciliation Best suited for generating separate keys between dealer and users.
VIII Conclusion
@@
Looking ahead, the insights gained from this work highlight several promising directions for future research. Exploring alternative multipartite states with configurations that allow users to access multiple modes could further enhance key rates and robustness in quantum communication protocols. Extending the protocol to networks with more than three users and examining the performance benefits in larger, more complex quantum networks is a natural progression. Developing advanced methods for channel parameter estimation that leverage access to multiple modes per user could further mitigate finite-size effects and improve the security and efficiency of the protocols. A crucial next step for this protocol is its experimental implementation, which presents a set of engineering challenges. For scalable, time-multiplexed resources, a central challenge is the engineering of high-speed optical switches to demultiplex and distribute modes to multiple users. While minimizing loss is ideal, the security framework can accommodate these losses if the device is trusted and well-characterized, shifting the problem towards precise calibration. Additionally, extending the security analysis to consider more sophisticated eavesdropping strategies would verify the protocol’s robustness against a wider range of attacks.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Acknowledgements
I would like to acknowledge Vladyslav Usenko, Ivan Derkach, Tobias Gehring, Christoph Pacher, Florian Kanitschar, and Ulrik Lund Andersen for their valuable discussions. A. O. acknowledges the project 8C22002 (CVStar) of the Czech Ministry of Education, Youth and Sports, which has received funding from the European Union’s Horizon 2020 research and innovation framework programme under grant agreement No. 731473 and 101017733, and the project No. 21-44815L of the Czech Science Foundation. A. O. also acknowledges the project IGA-PrF-2025-010 of Palacký University Olomouc.
@@ Appendix A Dual rail cluster state to six mode graph state
In Sec. V, we introduced a method to generate a pure six-mode graph state from a dual-rail cluster state. Here, we provide a justification for this method. Since we are working with Gaussian states, it suffices to show that, after appropriately grouping the nodes as suggested by the coloring scheme in Fig. 4(a), the covariance matrix of the resulting state corresponds to that of a pure six-mode graph state.
The dual-rail cluster state, denoted here by $\rho_{\text{dual}}$, consists of two parallel chains (rails) of quantum states: Rail $A$ and Rail $B$. Each node (quantum state) in these rails is uniquely identified by its rail ($A$ or $B$) and its position $i$ along that rail as illustrated in Fig. 10.
@@
Figure 10: Dual-rail cluster state $\rho_{\text{dual}}$ comprising two parallel rails (AA and $B$). Each node is labelled by its rail and position ii.
@@
where $\operatorname{Tr}^{(R)}_{i}$ denotes the partial trace over all modes except the one at position $i$ in rail $R$. We then form statistical mixtures by averaging the constituent states within each group. For each group $j\in\{1,2,3\}$ in rail $R$, the mixed state $\bm{\rho}^{(R)}_{j}$ is defined as:
$\bm{\rho}^{(R)}_{j}=\frac{1}{n}\sum_{m=1}^{n}\rho^{(R)}_{3m-(3-j)},$ (34)
@@
By forming these statistical mixtures, we create six mixed states in total (three groups in each of the two rails). The quadrature operators associated with these mixed states are denoted $\bm{x}^{(R)}_{j}$ and $\bm{p}^{(R)}_{j}$, where $j\in\{1,2,3\}$ and $R\in\{A,B\}$.
@@
$\left\langle\left(\bm{X}^{(R)}_{j}\right)^{2}\right\rangle=\frac{1}{n}\sum_{m=1}^{n}\left\langle\left(X^{(R)}_{3m-(3-j)}\right)^{2}\right\rangle,$ (35)
$\frac{1}{2}\left\langle[\bm{X}^{(R)}_{p},\bm{X}^{(S)}_{q}]_{+}\right\rangle=\frac{1}{2n}\sum_{m=1}^{n}\left\langle[X^{(R)}_{3m-(3-p)},X^{(S)}_{3m-(3-q)}]_{+}\right\rangle,$ (36)
@@
[…]
@@
$\mathbf{X}=\bigl($ $\bm{x}^{(A)}_{1},\bm{p}^{(A)}_{1},\bm{x}^{(B)}_{1},\bm{p}^{(B)}_{1},\bm{x}^{(A)}_{2},\bm{p}^{(A)}_{2}$, (38)
$\bm{x}^{(B)}_{2},\bm{p}^{(B)}_{2},\bm{x}^{(A)}_{3},\bm{p}^{(A)}_{3},\bm{x}^{(B)}_{3},\bm{p}^{(B)}_{3}\bigr)^{\!T}.$
With this ordering, the covariance matrix $\bm{\sigma}$ of the six-mode graph state is given by:
$\bm{\sigma}=\begin{pmatrix}\bm{V}&\bm{0}&\bm{C}&-\bm{C}&\frac{n-1}{n}\bm{C}&\frac{n-1}{n}\bm{C}\\ \bm{0}&\bm{V}&\bm{C}&-\bm{C}&-\frac{n-1}{n}\bm{C}&-\frac{n-1}{n}\bm{C}\\ \bm{C}&\bm{C}&\bm{V}&\bm{0}&\bm{C}&-\bm{C}\\ -\bm{C}&-\bm{C}&\bm{0}&\bm{V}&\bm{C}&-\bm{C}\\ \frac{n-1}{n}\bm{C}&-\frac{n-1}{n}\bm{C}&\bm{C}&\bm{C}&\bm{V}&\bm{0}\\ \frac{n-1}{n}\bm{C}&-\frac{n-1}{n}\bm{C}&-\bm{C}&-\bm{C}&\bm{0}&\bm{V}\end{pmatrix}$. (39)
Here, $\bm{V}$ and $\bm{C}$ are $2\times 2$ matrices representing the variances and covariances of the quadrature operators, respectively. These can be expressed as a function related to the squeezed quadrature variance $v$ and excess AS noise of the squeezed state employed in generating the dual rail cluster state Eq. 32. Specifically, $\bm{V}$ is the covariance matrix of a single mixed state, and $\bm{C}$ captures the correlations between different mixed states.
The factor $\frac{n-1}{n}$ within the covariance matrix is due to the fact that the count of vertices in a group that includes the extreme vertices cannot match the number of edges linking them. This factor makes the multipartite Gaussian state impure. One can purify this impurity, introduced due to finite states, by introducing additional modes. Alternatively, one can consider a pure six mode covariance matrix by removing the factor $\frac{n-1}{n}$. Introducing additional modes to purify the covariance matrix Eq. 39 would give us a tighter bound than considering the covariance matrix Eq. 30. That is, considering the pure six mode covariance matrix to estimate the Holevo bound is more pessimistic than considering a purified state corresponding to the covariance matrix Eq. 39 to estimate the Holevo bound. For sufficiently large $n$, the covariance matrix $\bm{\sigma}$ can be approximated to Eq. 30 a pure six mode state, and hence the improvement achieved through purification is insignificant.
@@ Appendix B Congruence Transformation for Covariance Matrices
Given an original covariance matrix $\bm{\Sigma}$ of dimensions $12n\times 12n$, of the modes of the dual-rail cluster state under consideration for grouping, we can derive the covariance matrix $\bm{\sigma}$ of dimensions $12\times 12$ corresponding to the six-mode graph state. This is achieved by applying a congruence transformation using a matrix $\mathbf{A}$ of size $12\times 12n$, such that:
$\bm{\sigma}=\frac{1}{n}\mathbf{A}\bm{\Sigma}\mathbf{A}^{\intercal}.$ (40)
@@ B.1 Congruence transformation for adjacency matrices
A complex-weighted adjacency matrix is used in continuous-variable quantum computing to represent Gaussian pure states, extending the real-weighted graphs of ideal cluster states to more realistic, physical states [40]. This formalism simplifies the analysis of entanglement structures, Gaussian operations, and deviations from ideal behaviors, which are essential for applications in CV quantum computing. Thus, we define a similar transformation to the one used in the previous section, enabling us to derive the adjacency matrix for a six-mode graph state $\bm{Z}_{six-mode}$ from the adjacency matrix of a dual-rail cluster state $\bm{Z}_{dual}$.
$\bm{Z}_{six-mode}=\frac{1}{n}\bm{B}\bm{Z}_{dual}\bm{B^{T}}$ (43)
@@
(b) Complex weighted graph of quotient graph derived from dual rail cluster state.
@@
Appendix C Rationale for the coloring strategy
To convert the infinitely repeating dual-rail cluster lattice into a finite resource that is easy to manipulate in security proofs, we map groups of physical modes onto color classes and then merge every class into a single vertex; the resulting state is called a quotient graph state. The coloring we adopt is engineered to satisfy two operational requirements:
  1. (i)
    Neighbourhood preservation — The quotient must retain all edges that enter the closed neighbourhood $N[v]$ of at least one vertex $v$ in the original lattice, so that any pair of modes that were immediate neighbours remain directly correlated after the homomorphism.
  2. (ii)
    Purity —The quotient state should be a pure Gaussian graph state; purity simplifies both covariance-matrix algebra and the security analysis in Sec. III.
The procedure is generic and can be reused for other cluster topologies; we describe it once here and apply it to the dual-rail lattice below.
Step 1: Identify a unit cell that encloses a full closed neighbourhood.
Let $G=(V,E)$ be the graph of the open-ended lattice. We search for the smallest induced subgraph $G_{\text{cell}}\subset G$ that contains a vertex $v$ together with all vertices in $N[v]$. For the dual-rail lattice, this unit cell contains six vertices: two from each rail at positions 3m23m-2, 3m13m-1, and $3m$ ($m\in\mathbb{Z}$).
Step 2: Assign one color per vertex inside the cell.
Each vertex of $G_{\text{cell}}$ receives a distinct color $c\in{1,\dots,|V(G_{\text{cell}})|}$. Translational tiling of the cell propagates this coloring across the infinite lattice, producing exactly six global color classes in the dual-rail case.
Step3: Form the quotient graph.
Vertices that share the same color are identified (fused) into a single vertex. If colors $c_{i}$ and $c_{j}$ appeared on adjacent physical vertices, the corresponding quotient vertices are connected by an edge. Algebraically, if $Z_{\text{dual}}$ is the complex-weighted adjacency matrix of the dual-rail lattice, the quotient adjacency matrix is obtained via the congruence transformation.
$Z_{\text{quot}}=\frac{1}{n}BZ_{\text{dual}}B^{\mathsf{T}},$ (45)
where the selector matrix $B$ is defined in Sec. B and $n$ is the number of cell repetitions.
Step 4: Verify purity and neighbourhood fidelity.
The same transformation applied to the lattice covariance matrix $\Sigma$ yields
$\sigma=\frac{1}{n}A\Sigma A^{\mathsf{T}},$ (46)
with $A$ given in Eq. 42. Taking the limit $n\to\infty$ removes boundary terms proportional to (n1)/n(n-1)/n, leaving the six-mode covariance matrix of Eq. 30 which is pure ($Det[\sigma]=1$). Moreover, every edge that linked a vertex to its first neighbours inside $G_{\text{cell}}$ survives in $Z_{\text{quot}}$, so immediate–neighbour correlations are identical to those of the parent lattice.
Result for the dual-rail lattice. Applying the above steps to the dual-rail cluster yields the coloring shown in Fig. 4(a). The six color classes correspond to the vertices $a_{1},a_{2},b_{1},b_{2},c_{1},c_{2}$} of Fig. 4(b). This six-mode pure quotient graph state is the smallest resource that simultaneously preserves all closed–neighbourhood correlations and admits a compact security proof.
Reuse for other lattices. For any future cluster geometry one simply (i) locates a cell containing a full neighbourhood, (ii) assigns unique colors to its vertices, and (iii) repeats the tiling. The size of the quotient graph is bounded below by the cardinality of that neighbourhood and above by the chromatic number of the lattice.
Appendix D Parameter estimation
@@
Here, $x^{(i)}$ denotes a squeezed quadrature with variance $V$, while $p^{(i)}$ denotes the corresponding anti-squeezed quadrature with variance $V_{N}+\frac{1}{V}$.
@@
$\mathrm{Var}[x_{d}\tilde{x}_{u}]$ $=\tfrac{t}{16}\Bigl(5\bigl(V_{N}+\tfrac{1}{V}\bigr)^{2}+5V^{2}+6\bigl(V_{N}+\tfrac{1}{V}\bigr)V\Bigr)$ (63)
$\quad+(1-t)\left(\tfrac{1+V(V_{N}+V)}{2V}\right).$
@@
$x_{1}$ $=\frac{\sqrt{2}\,p^{(1)}+x^{(2)}+\sqrt{3}\,x^{(3)}}{\sqrt{6}},$ (64)
$x_{2}$ $=\frac{\sqrt{2}\,x^{(2)}-p^{(1)}}{\sqrt{3}},$
$x_{3}$ $=\frac{-\sqrt{2}\,p^{(1)}-x^{(2)}+\sqrt{3}\,x^{(3)}}{\sqrt{6}}.$
@@
$\mathrm{Var}[x_{d}\tilde{x}_{u}]$ $=\tfrac{t}{9}\Bigl(2\bigl(V_{N}+\tfrac{1}{V}\bigr)^{2}+5V^{2}+2\bigl(V_{N}+\tfrac{1}{V}\bigr)V\Bigr)$ (65)
$\quad+(1-t)\left(\tfrac{1+V(V_{N}+2V)}{3V}\right).$
@@
$x_{d}=\frac{x^{(1)}+p^{(2)}}{\sqrt{2}},\quad x_{u}=\frac{1}{2}(\sqrt{2}\,p^{(0)}-x^{(1)}+p^{(2)}).$ (66)
@@
$\mathrm{Var}[x_{d}\tilde{x}_{u}]$ $=\tfrac{t}{8}\left(V_{N}+\tfrac{1}{V}\right)\left(V_{N}+\tfrac{1}{V}+1\right)$ (67)
$\quad+V(1+V)$
$\quad+(1-t)\left(\tfrac{1+V(V_{N}+V)}{2V}\right).$